# Proprioceptive Image: An Image Representation of Proprioceptive Data from Quadruped Robots for Contact Estimation Learning

### v1

@@ III-B Simulation Results
To further evaluate our PI formulation, we generated a dataset in the MuJoCo simulator [31] using a 60kg Unitree B2 quadruped robot. The dataset includes three environments (flat, uneven, and pyramid), two gaits (trot and crawl), and both stable ($\mu=1.0$) and slippery ($\mu=0.2$) terrain conditions. The uneven and pyramid environments were randomized in relation to the boxes positions and the height of the pyramid. The robot traversed these environments under a gradient-based MPC controller developed in [32], with trunk linear velocities ranging from –1 to 1 m/s and angular velocities from –0.3 to 0.3 rad/s. Proprioceptive signals were recorded, including joint positions, joint velocities, foot positions, foot velocities, IMU signals and GRFs with Gaussian noise added to better approximate real-world sensing conditions. Contact estimation labels were directly extracted from the simulator and used as ground truth for training and evaluation. In total, the dataset comprises more than 1.5M synced data, organized by robot gait and environment stability. Each subset was further divided into training, validation, and test splits to enable fair comparison between our PI formulation, the approach of [24], and a baseline method based on GRF threshold [6, 18]. Figure 9 illustrates examples of scenes from the simulated dataset across the different environments.
@@ IV Discussion
Although PI encodings provide rich learnable patterns, their creation is relatively slow (1.5 ms per image) due to the current Python implementation running on the CPU. We expect that a GPU-based implementation would substantially reduce generation time and make the formulation more practical for real-time robotic systems. The PI-based contact estimation model achieved a mean inference time of 8.5 ms on GPU, corresponding to a processing frequency of approximately 120 Hz.
@@ V CONCLUSIONS
This paper presented a novel formulation of proprioceptive signals into an image representation. Each image representation encodes four important temporal information of dynamic measurement behavior and are concatenated based on the quadruped robot morphology. The PI formulation was evaluated on the contact estimation problem, using simple CNN architectures trained on both simulated and real-world data, achieving the best results in comparison with state-of-the-art methods. The PI is the first step toward deeper representations of robot sensing. For future work, we expect that the PI can be integrated into more advanced learning structures such as Graph Neural Networks similar to [26] and also applied PIs as input for Reinforcement Learning techniques. Moreover, combining a PI with exteroceptive sensors, such as cameras, could yield a unified representation of both the robot’s internal state and its external environment, further enhancing perception and state learning. Finally, we aim to explore the applicability of PI in other robotic platforms, including humanoids and manipulators.

### v2

@@ III-B Simulation Results
To further evaluate our PI formulation, we generated a dataset in the MuJoCo simulator [23] using a 60kg Unitree B2 quadruped robot. The dataset includes three environments (flat, uneven, and pyramid), two gaits (trot and crawl), and both stable ($\mu$=1.0) and slippery ($\mu$=0.2) terrain conditions, where $\mu$ is the friction coefficient. The uneven and pyramid environments were randomized in relation to the boxes positions and the height of the pyramid. The robot traversed these environments under a gradient-based MPC controller developed in [24], with trunk linear velocities ranging from –1 to 1 m/s and angular velocities from –0.3 to 0.3 rad/s. Proprioceptive signals were recorded, including joint positions, joint velocities, foot positions, foot velocities, IMU signals and GRFs with Gaussian noise added to better approximate real-world sensing conditions. Contact estimation labels were directly extracted from the simulator and used as ground truth for training and evaluation. In total, the dataset comprises more than 1.5M synced data samples, organized by robot gait and environment stability. Each subset was further divided into training, validation, and test splits to enable fair comparison between our PI formulation, the approach of [12], and a baseline method based on GRF threshold [18, 7]. Figure 9 illustrates examples of scenes from the simulated dataset across the different environments.
@@ IV Discussion
Although PI encodings provide rich, learnable patterns, their generation is relatively slow (1.5 ms per image for $w=10$) due to the current CPU-based Python implementation. Increasing image resolution can provide richer feature content and improve accuracy, but it also increases encoding cost. In our experiments, the model achieved a mean GPU inference time of 8.5 ms ($\approx 120Hz$). We therefore expect that a multi-threaded C++ implementation will substantially reduce generation time, making the formulation more suitable for real-time robotic systems. Beyond runtime, PIs derived from IMU signals exhibit Brownian-walk-like drift, which over long experiments leads to saturated, repetitive patterns. Finally, the kernel functions used to encode each PI component were heuristically tuned for contact estimation, suggesting that a more systematic study is needed to generalize the encoding to other tasks and robotic platforms.
@@ V CONCLUSIONS
This paper presented a novel formulation of proprioceptive signals into an image representation. Each image representation encodes four important types of temporal information related to dynamic measurement behavior and is concatenated based on the morphology of the quadruped robot. The PI was evaluated on the contact estimation problem, using a simple CNN architecture trained on both simulated and real-world data, achieving the best results in comparison with state-of-the-art methods. The PI is the first step toward deeper representations of robot sensing. For future work, we expect that the PI can be integrated into other learning structures such as GNN similar to [3] and also applied as input for reinforcement learning techniques. Moreover, combining a PI with exteroceptive sensors, such as cameras, could yield a unified representation of both the robot’s internal state and its external environment, further enhancing perception and state learning. Finally, we aim to explore the applicability of PI in other robotic platforms, including humanoids and manipulators.